Read the t-distribution table and answer the following: For a two-tailed hypothesis testing with 36 degrees of freedom using LaTeX: \alphaα=0.05, the critical value (-/+) LaTeX: t_\frac{\alpha}{2}t α 2 = ____.
Answer #1

solution

Degrees of freedom = df =36

= 0.05

  / 2= 0.05 / 2 = 0.025

t /2,df = t0.025,36 =2.028 ( using student t table)